A rubber band has a spring constant of 45N\m. If the rubber band is pulled by a force of 9.8N, (a) how far would it stretch?(b)what is the worked done by this force in stretching this string

1
Expert's answer
2021-12-06T09:49:39-0500

(a)

"F=kx,""x=\dfrac{F}{k}=\dfrac{9.8\ N}{45\ \dfrac{N}{m}}=0.218\ m."

(b)

"W=\dfrac{1}{2}kx^2=\dfrac{1}{2}\times45\ \dfrac{N}{m}\times(0.218\ m)^2=1.07\ J."